Suppose an investor buys a share of Sony at a price of $38,720 at the start of the year. During the year, the investor receives a dividend of ¥500. At the end of the year, the price of Sony is ¥49,560. During the year, the exchange rate goes from $1 ¥150 to $1 -Y175. What was the investors dollar return on Sony (in percent)?
Answer #1

initial investment = 38,720/150

final value of investment = (500 + 49,560)/175 = 50,060/175

dollar return = (50,060/175)/(38,720/150) - 1

dollar return = 10.82%